OpenStudy (anonymous):

Please Help Me. Thank You (We have to solve for x) 1. 7/8 - 1/x = 3/4 2. 1/2 + 4/2x = x+4/10 3. 3/3x + 1/x+4 = 10/7x 4. 4/x + 4/x^2-9 = 3/x-3 5. 2/x-2 + 7/x^2-4 = 5/x

OpenStudy (unklerhaukus):

1. \[\frac{7}{8}-\frac1 x =\frac 3 4\] times everything by something that will make those denominators go away {some thing like 8x}

OpenStudy (anonymous):

I still don't understand it.

OpenStudy (anonymous):

Getting the least common denominator...multiplying both the denominator 8 ans x is a good strategy...then you can procees as fallows.. \[(7x-8)/8x = 3/4\]

OpenStudy (anonymous):

For #1 x= 4

OpenStudy (anonymous):

No it should be 8..

OpenStudy (anonymous):

For #2, I got x=-4 and x=5

OpenStudy (anonymous):

Could not be.,,x only has one value to satisfy the equation...

OpenStudy (anonymous):

Okay. Let me correct it.

OpenStudy (anonymous):

For #2 is it x=-4

OpenStudy (anonymous):

For #3 is x=3

OpenStudy (anonymous):

For #4 is it x=-4 x=9

OpenStudy (anonymous):

For #5 is it x= 4/3 x=5

OpenStudy (anonymous):

For #2 I guess the equation is quadratic...I simplify it to \[20x ^{2} -2x+40=0 \] use quadratic formula to solve for x...

OpenStudy (anonymous):

Okay what about the other ones

OpenStudy (anonymous):

Quadratic also...simply the equation...use the quadratic formula...or factoring method..to solve for x...so x may have two different values...

OpenStudy (anonymous):

When i used the quadratic formula i got x=-3 x=-1 .. (Now im confuse)

OpenStudy (anonymous):

for # 2 I guess x turns out to be imaginary number... By the..can u check for #5 I think x = 5, x = -4/3

OpenStudy (anonymous):

For #2 it would be x=-4

OpenStudy (anonymous):

always check your answer by substituting the value of x to the original equation...for #2\[1/2 + 4/2(-4) = -4 + 4/10\] ---> 1/2 +(-4/8) = (-40+4)/10 0=-36/10 ....since the right side of the equation does not equal to the left side..then x =-4 must not be the answer right?

OpenStudy (anonymous):

I tried the answer 5 but something doesn't look right. 1/2 + 4/ 2(5) = (5)+4/10 I got 1/2 + 4/10= 9/10

OpenStudy (anonymous):

Can u verify for # 2 equation...\[1/2 + 4/2x = x + (4/10) \] I guess the equation is like this above.. not ..\[1/2 + 4/ 4x = (x+4) /10\]

OpenStudy (anonymous):

It's 1:17. I need to submit this questions. I need to move on to another subject. Thanks for your help. :(

OpenStudy (anonymous):

Ok..gud luck..!